3. Finding the optimal combination of two goods Nick enjoys consuming both guacamole and salsa. Each jar of guacamole costs Pc $1, and each jar of salsa costs Ps $2. Suppose that Nick buys 75 jars of guacamole and 50 jars of salsa per month. The following graphs show his marginal utility curves for guacamole and salsa. At his current consumption level, Nicks marginal utility from consuming the last jar of guacamole he bought is MUG 12 utils per jar, and his marginal utility from consuming the last jar of salsa he bought is MUs 12 utils per jar. GUACAMOLE (Jars) SALSA (Jars)
Answer #1

Nick's utility is maximised when marginal utility per dollar is same for both goods. Given , his current consumption , Nick's marginal utility per dollar for guacamole = 12/1 = 12. And marginal utility per dollar for Salsa = 12/2 = 6.

These two are not the same. It implies that Nick is not maximising his utility at this bundle .And because salsa costs more than guacamole. therefore, Nick's utilty can maximised if he spend less on salsa and more of Guacamole.

Hence, option(D) is correct. No, Nick's is not maximise his current utility. He could buy more guacamole and less salsa , not spend any more money and be better off.

Guacamole Salsa
Combination Jars MUG/PG Jars MUS/PS Consumer equilibrium
A 100 8/1 =8 100 8/2=4 NO
B 75 12/1=12 50 12/2=6 NO
C 125 4/1 =4 100 8/2=4 YES
D 25 20/1=20 25 14/2=7 NO

Suppose Nick has a budget of $325 per moneth , which he spends on Guacomole and salsa. Yo maximise his utility , Nick will purchase 125 jars of Guacamole and 100 jars of salsa per month. Because then MUG/PG= MUS/PS =4.